OpenStudy (anonymous):

You have a 5-question multiple-choice test. Each question has four choices. You don’t know any of the answers. What is the experimental probability that you will guess exactly three out of five questions correctly?

OpenStudy (anonymous):

Hmm good question.

OpenStudy (anonymous):

@mathteacher1729 this looks like a question for you

OpenStudy (kropot72):

The number of combinations of questions taken 3 at a time is\[\left(\begin{matrix}5 \\ 3\end{matrix}\right)=\frac{5!}{2!\times 3!}=10\] The probability of each combination of 3 questions having all correct answers is \[(\frac{1}{4})^{3}=0.015625\] Therefore the probability of quessing exactly 3 out of the five questions correctly is 0.015625*10=? Can you finish it?

OpenStudy (anonymous):

dont get it

OpenStudy (kropot72):

Which part don't you understand?